# Large-scale distributed synchronization systems, using a cancel-on-completion redundancy mechanism

### v3

@@ 1.1 Model introduction and motivation
One can view the process described above as a particle system, if each server is associated with a “particle,” and its workload is interpreted as the particle “location” on the real line. With this view, an arriving class-$j$ “job” selects $d_{j}$ particles uniformly at random, and their random non-negative potential jump sizes are drawn from distribution $F_{j}$, which determine their potential new locations after the jumps. However, the actual new locations the selected particles jump to, are “truncated” by the $k_{j}$-th smallest potential new location (due to the c.o.c. mechanism). Between “job” arrivals, each particle moves to the “left” at constant speed $v=1$, but cannot move to left of the “regulation” boundary at 0; a particle system with this feature will be referred to as left-regulated. Due to the jump truncation feature (c.o.c. mechanism) the system is “non-work-conserving” in the sense that the expected total displacement of the selected particles after a class-$j$ job arrival is not constant – it, in general, depends on the system state at the time of arrival, namely on the relative locations of selected particles. The non-work-conservation is the model feature which substantially complicates the system analysis (see section 1.1 in [15]).
@@
To motivate our more general setting, consider the following example of a multi-agent distributed synchronization model. We have $n$ agents (particles), each trying to advance its own state (particle location), but there are dependencies between the agents. (One specific motivation for it is a parallel simulation, where an agent is a processor, simulating an element, or a subsystem, of a larger physical system. An agent state is the local simulation time of the processor. Cf. [11].) A job arrival corresponds to selection of a groups of agents, of size $d$, that are currently interdependent. Upon such selection, the agents in the group try to independently advance their states (locations) by random distances (jump sizes) $(\xi_{1},\ldots,\xi_{d})$, but after they advance, they have to synchronize – roll back to the left-most new location within the group. (In terms of parallel simulation, the jump size of an agent corresponds to the time duration by which the agent can advance its local simulation without affecting other agents in the group.) If we assume that jobs arrive as a Poisson process, and the groups are selected uniformly at random, we obtain a single-class unregulated (free) system with $(d,1)$-c.o.c. redundancy.
Results of this paper, in particular, characterize the limits $\lim_{n\to\infty}v_{n}$ of the average velocity $v_{n}$ for general free systems. This characterization is particularly sharp in the special case when the component sizes $(\xi^{(j)}_{1},\ldots,\xi^{(j)}_{d_{j}})$ for each class are independent identically distributed, with a distribution having increasing hazard rate.
@@ 3.1 Multi-server queueing system with cancel-on-completion redundancy
Consider a queueing model with $n$ identical servers, each processing its work at rate $1$. The workload $w$ of a server at a given time $t$ is its unfinished work, namely the time duration until the server becomes idle assuming no new job arrivals after $t$. There is a finite set $\cal J$} of job classes $j$. Jobs arrive as Poisson process of rate $\lambda^{n}n$, for some $\lambda^{n}\geq 0$. The probability that an arriving job is of class $j$ is $\pi_{j}^{n}\geq 0$, $\sum_{j}\pi_{j}^{n}=1$. We will also use notation $\sigma_{j}^{n}=\lambda^{n}\pi_{j}^{n}$ for the class $j$ arrival rate per server.
Each job class $j$ has three parameters: integers $k_{j}$ and $d_{j}$ (they will not depend on $n$) such that $1\leq k_{j}\leq d_{j}$, and the exchangeable probability distribution $F_{j}$ on $\mathbb{R}_{+}^{d_{j}}$. (Exchangeability of $F_{j}$ means that it is invariant w.r.t. permutations of components.) When a class $j$ job arrives, $d_{j}$ servers are selected uniformly at random (without replacement); these servers form the selection set of the job. The job places $d_{j}$ components on the selected servers; the component sizes (potential amounts of required service) $(\xi^{(j)}_{1},\ldots,\xi^{(j)}_{d_{j}})$ are drawn according to distribution $F_{j}^{n}$, independently of the process history up to the job arrival time. Each server processes its work (components of different jobs) in the First-Come-First-Served (FCFS) order. A job of class $j$, to be completed, requires $k_{j}$ (out of $d_{j}$) components to be processed, and as soon as $k_{j}$ components of the job complete their service (i.e. receive the amounts of service equal to their sizes), the unfinished service of the remaining $d_{j}-k_{j}$ components of the job is “canceled” and the job immediately leaves the system; we will call this $(d_{j},k_{j})$-c.o.c. redundancy. Throughout the paper we use notation $\bar{d}\doteq\max_{j}d_{j}$.
@@ 3.3 Assumptions on job component size distributions
For any $j$, $\mathbb{E}[\xi_{1}^{(j)}]^{2}<\infty$ (and then $\mathbb{E}[\sum_{i}\xi_{i}^{(j)}]^{2}<\infty$).
@@ 3.5 Steady-state asymptotic independence definition
For the anchored system with frame $0,\infty)$ (i.e. left-regulated at 0 and unregulated on the right), the following property will be called SSAI for the full range of loads (SSAI-FRL): there exists $\underline{v}$ such that for any $v>\underline{v}$ the convergence $x^{n}(\infty)\Rightarrow x^{*,v}$ holds for some fixed $x^{*,v}\in\tilde{\cal X}^{pr}$ and $x^{*,v}_{0}\uparrow 1$ as $v\downarrow\underline{v}$. The name SSAI-FRL is natural, because in terms of the multi-server queueing system $x^{*,v}_{0}$ is the limiting fraction of occupied servers – the system load. The notion of SSAI-FRL was introduced in [15]; see [15, section 1.2] for a discussion of this property is not automatic – and may be challenging to prove – for non-work-concerving systems.
@@ 4.1 Basic monotonicity and continuity
Consider a fixed set of particles, selected by a class-$j$ job, and labeled WLOG by $1,2,\ldots,d_{j}$, with (deterministic) locations given by vector $W=(W_{1},\ldots,W_{d_{j}})$. Let $\xi^{(j)}=(\xi^{(j)}_{1},\ldots,\xi^{(j)}_{d_{j}})$ be the (deterministic) vector of their jump sizes. Denote by $\hat{W}=(\hat{W}_{1},\ldots,\hat{W}_{d_{j}})$ the vector of the particles’ locations after the job arrival. Then $\hat{W}$ is monotone non-decreasing and continuous in both $W$ and $\xi^{(j)}$.
This fact easily implies properties like: two versions of the same process, one starting from a smaller state than another, can be coupled so that the state of the former is smaller than the state of the latter at all times. This property obviously prevails, for example, if, in addition, the distribution of jump sizes in the former process is dominated by that in the latter, and/or if movement of particles in the former system in “impeded” by a fixed or moving right boundary, and/or if movement of particles in the latter system in “helped” by a fixed or moving left boundary, and so on. Similarly, for example, if two processes start from the same initial state and the distributions of component sizes in the former converges to those in the latter, then the former process converges (in appropriate sense) to the latter; analogous continuity properties hold as well. Furthermore, such monotonicity and continuity properties are naturally inherited by the mean-field limits, which we introduce in Section 5. All such properties are straightforward – we will not spell out all of them. (Some are spelled out in section 6 of [15].) In the rest of this paper, we simply refer to them, where necessary, as monotonicity and/or continuity.
@@ 6 Preliminary definitions and results (generalizing those in [15]), related to SSAI
Consider a free system. Suppose Assumptions 3.3, 3.4, and 6.2 hold. Then, there exists $\bar{n}$ such that for all $n\geq\bar{n}$, $\mathring{x}^{n}(\cdot)$ is positive recurrent. [^1] Assumption  3.4 is not needed for positive recurrence. See Theorem A.1 in Appendix A. If, in addition, Assumption 3.2 holds, then there exist $\bar{C}>0$ such that
@@ 8 Proofs
$\limsup_{t\downarrow 0}\sup_{w}\frac{P^{t}-I}{t}[(w^{2}\wedge C_{2})-A_{C_{2}}(w)]\leq 0.$
@@
$0=\lim_{t\downarrow 0}\mathbb{E}\left[(X(t)^{2}\wedge C_{2})-(X(0)^{2}\wedge C_{2})\right]\leq\mathbb{E}A_{C_{2}}(X(\infty))=\mathbb{E}[2X(\infty)\bar{\psi}_{X(\infty)}+C]{\bf I}\{X(\infty)^{2}\leq C_{2}\},$ (40)
@@
If there is no left boundary, $\bar{\psi}_{w}$ is continuous, and there is a point $u$ such that $\bar{\psi}_{u}=0$. Let us choose this point for centering and consider Lyapunov function $(w-u)^{2}\wedge C_{2}$. Using analogous generator upper bound, we can show that $\mathbb{E}|X(\infty)-u|=\infty$ is impossible, because it would imply negative average drift of $X(t)-u]^{2}\wedge C_{2}$ when $C_{2}$ is large. Therefore, $\mathbb{E}|X(\infty)-u|<\infty$, and $\bar{x}$ is well-defined finite. (A small subtlety here. If $X(t)$ jumps so that $X(t+)>u+\sqrt{C_{2}}$, then replacing $X(t+)-u]^{2}\wedge C_{2}=C_{2}$ by $X(t+)-u]^{2}$ in the drift estimate, makes the estimate larger.)
@@
$0\leq\mathbb{E}[2X^{(\kappa)}(\infty)\bar{\psi}_{X^{(\kappa)}(\infty)}+C]{\bf I}\{X^{(\kappa)}(\infty)^{2}\leq C_{2}\}.$

### v4

@@ 1.1 Model introduction and motivation
One can view the process described above as a particle system, if each server $i$ is associated with a “particle,” and its workload $W_{i}$ is interpreted as the particle “location” on the real line. With this view, an arriving class-$j$ “job” selects $d_{j}$ particles uniformly at random, and their random non-negative potential jump sizes $(\xi^{(j)}_{1},\ldots,\xi^{(j)}_{d_{j}})$ are drawn from distribution $F_{j}$. So, if a selected particle $i$ current location is $W_{i}$, its potential new location is $W_{i}+\xi_{i}^{(j)}$. However, the actual new location particle $i$ jumps to is “truncated” at point $W^{*}$, which is the $k_{j}$-th smallest potential new location among the selected particles (due to the c.o.c. mechanism), so that the particle actually jumps to $\max\{W_{i},\min\{W_{i}+\xi_{i}^{(j)},W^{*}\}\}$. Between “job” arrivals, each particle moves to the “left” at constant speed $v=1$, but cannot move to the left of the “regulation” boundary at 0; a particle system with this feature will be referred to as left-regulated. Due to the jump truncation feature (c.o.c. mechanism) the system is “non-work-conserving” in the sense that the expected total displacement of the selected particles after a class-$j$ job arrival is not constant – it, in general, depends on the system state at the time of arrival, namely on the relative locations of selected particles. The non-work-conservation is the model feature which substantially complicates the system analysis (see section 1.1 in [15]).
@@
To motivate our more general setting, consider the following example of a multi-agent distributed synchronization model. We have $n$ agents (particles), each trying to advance its own state (particle location), but there are dependencies between the agents. (One specific motivation is a parallel simulation, where an agent is a processor, simulating an element, or a subsystem, of a larger physical system. An agent state is the local simulation time of the processor. Cf. [11].) A job arrival corresponds to selection of a group of agents, of size $d$, that are currently interdependent. Upon such selection, the agents in the group, first, separately advance their states (locations) $W_{1},\ldots,W_{d}$ forward by random distances (potential jump sizes) $\xi_{1},\ldots,\xi_{d}$, but then have to synchronize – each agent $i$ rolls back its new state from $W_{i}+\xi_{i}$ to $\max\{W_{i},\min_{\ell}(W_{\ell}+\xi_{\ell})\}$. (In terms of parallel simulation, $\xi_{i}$ is the time duration by which processor $i$ can advance its local simulation without affecting simulations run by other processors in the group.) If we assume that jobs arrive as a Poisson process, and the groups are selected uniformly at random, we obtain a single-class unregulated (free) system with $(d,1)$-c.o.c. redundancy.
Results of this paper, in particular, characterize the limits $\lim_{n\to\infty}v_{n}$ of the average velocity $v_{n}$ for general free systems. This characterization is particularly sharp in the special case when the component sizes $(\xi^{(j)}_{1},\ldots,\xi^{(j)}_{d_{j}})$ for each class are independent identically distributed, with the distribution having increasing hazard rate.
@@ 3.1 Multi-server queueing system with cancel-on-completion redundancy
Consider a queueing model with $n$ identical servers, each processing its work at rate $1$. The workload of a server at a given time $t$ is its unfinished work, namely the time duration until the server becomes idle assuming no new job arrivals after $t$. There is a finite set $\cal J$} of job classes $j$. Jobs arrive as Poisson process of rate $\lambda^{n}n$, for some $\lambda^{n}\geq 0$. The probability that an arriving job is of class $j$ is $\pi_{j}^{n}\geq 0$, $\sum_{j}\pi_{j}^{n}=1$. We will also use notation $\sigma_{j}^{n}=\lambda^{n}\pi_{j}^{n}$ for the class $j$ arrival rate per server.
Each job class $j$ has three parameters: integers $k_{j}$ and $d_{j}$ such that $1\leq k_{j}\leq d_{j}$, and an exchangeable probability distribution $F_{j}$ on $\mathbb{R}_{+}^{d_{j}}$. (Exchangeability of $F_{j}$ means that it is invariant w.r.t. permutations of components.) When a class $j$ job arrives, $d_{j}$ servers are selected uniformly at random (without replacement); these servers form the selection set of the job. The job places $d_{j}$ components on the selected servers; the component sizes (potential amounts of required service) $(\xi^{(j)}_{1},\ldots,\xi^{(j)}_{d_{j}})$ are drawn according to distribution $F_{j}$, independently of the process history up to the job arrival time. Each server processes its work (components of different jobs) in the First-Come-First-Served (FCFS) order. A job of class $j$, to be completed, requires $k_{j}$ (out of $d_{j}$) components to be processed, and as soon as $k_{j}$ components of the job complete their service (i.e. receive the amounts of service equal to their sizes), the unfinished service of the remaining $d_{j}-k_{j}$ components of the job is “canceled” and the job immediately leaves the system; we will call this $(d_{j},k_{j})$-c.o.c. redundancy. Throughout the paper we use notation $\bar{d}\doteq\max_{j}d_{j}$.
@@ 3.3 Assumptions on job component size distributions
For any $j$, $\mathbb{E}(\xi_{1}^{(j)})^{2}<\infty$ (and then $\mathbb{E}(\sum_{i}\xi_{i}^{(j)})^{2}<\infty$).
@@ 3.5 Steady-state asymptotic independence definition
For the anchored system with frame $0,\infty)$ (i.e. left-regulated at 0 and unregulated on the right), the following property will be called SSAI for the full range of loads (SSAI-FRL): there exists $\underline{v}$ such that for any $v>\underline{v}$ the convergence $x^{n}(\infty)\Rightarrow x^{*,v}$ holds for some fixed $x^{*,v}\in\tilde{\cal X}^{pr}$ and $x^{*,v}_{0}\uparrow 1$ as $v\downarrow\underline{v}$. The name SSAI-FRL is natural, because, in terms of the multi-server queueing system, $x^{*,v}_{0}$ is the limiting fraction of occupied servers – the system load. The notion of SSAI-FRL was introduced in [15]; see [15, section 1.1] for a discussion of this property, which is not automatic – and may be challenging to prove – for non-work-concerving systems.
@@ 4.1 Basic monotonicity and continuity
Consider a fixed set of particles, selected by a class-$j$ job, and labeled WLOG by $1,2,\ldots,d_{j}$, with (deterministic) locations given by vector $W=(W_{1},\ldots,W_{d_{j}})$. Let $\xi^{(j)}=(\xi^{(j)}_{1},\ldots,\xi^{(j)}_{d_{j}})$ be the (deterministic) vector of their potential jump sizes. Denote by $\hat{W}=(\hat{W}_{1},\ldots,\hat{W}_{d_{j}})$ the vector of the particles’ locations after the job arrival. Then $\hat{W}$ is monotone non-decreasing and continuous in both $W$ and $\xi^{(j)}$.
This fact easily implies properties like: two versions of the same process, one starting from a smaller state than another, can be coupled so that the state of the former is smaller than the state of the latter at all times. This property obviously prevails, for example, if, in addition, the distribution of potential jump sizes in the former process is dominated by that in the latter, and/or if movement of particles in the former system is “impeded” by a fixed or moving right boundary, and/or if movement of particles in the latter system is “helped” by a fixed or moving left boundary, and so on. Similarly, for example, if two processes start from the same initial state and the distribution of potential jump sizes in the former converges to those in the latter, then the former process converges (in appropriate sense) to the latter; analogous continuity properties hold as well. Furthermore, such monotonicity and continuity properties are naturally inherited by the mean-field limits, which we introduce in Section 5. All such properties are straightforward – we will not spell out all of them. (Some are spelled out in section 6 of [15].) In the rest of this paper, we simply refer to them, where necessary, as monotonicity and/or continuity.
@@ 6 Preliminary definitions and results (generalizing those in [15]), related to SSAI
Consider a free system. Suppose Assumptions 3.3, 3.4, and 6.2 hold. Then, there exists $\bar{n}$ such that for all $n\geq\bar{n}$, $\mathring{x}^{n}(\cdot)$ is positive recurrent. [^1] In fact, Assumption  3.4 is not needed for positive recurrence. See Theorem A.1 in Appendix A. If, in addition, Assumption 3.2 holds, then there exist $\bar{C}>0$ such that
@@ 8 Proofs
$\limsup_{t\downarrow 0}\sup_{w}\left[\frac{P^{t}-I}{t}(w^{2}\wedge C_{2})-A_{C_{2}}(w)\right]\leq 0.$
@@
$0=\lim_{t\downarrow 0}\mathbb{E}\frac{(X(t)^{2}\wedge C_{2})-(X(0)^{2}\wedge C_{2})}{t}\leq\mathbb{E}A_{C_{2}}(X(\infty))=\mathbb{E}\left[[2X(\infty)\bar{\psi}_{X(\infty)}+C]{\bf I}\{X(\infty)^{2}\leq C_{2}\}\right],$ (40)
@@
If there is no left boundary, $\bar{\psi}_{w}$ is continuous, and there is a point $u$ such that $\bar{\psi}_{u}=0$. Let us choose this point for centering and consider Lyapunov function $(w-u)^{2}\wedge C_{2}$. Using analogous generator upper bound, we can show that $\mathbb{E}|X(\infty)-u|=\infty$ is impossible, because it would imply negative average drift of $X(t)-u]^{2}\wedge C_{2}$ when $C_{2}$ is large. Therefore, $\mathbb{E}|X(\infty)-u|<\infty$, and $\bar{x}$ is well-defined finite. (A small subtlety here. If $X(t)$ jumps so that $X(t+)>u+\sqrt{C_{2}}$, then replacing $X(t+)-u]^{2}\wedge C_{2}=C_{2}$ by $X(t+)-u]^{2}$ in the drift estimate makes the estimate larger.)
@@
$0\leq\mathbb{E}\left[[2X^{(\kappa)}(\infty)\bar{\psi}_{X^{(\kappa)}(\infty)}+C]{\bf I}\{X^{(\kappa)}(\infty)^{2}\leq C_{2}\}\right].$